A garden plot 5m by 15m has one of its longest sides next to a wall. The area of the plot is to be doubled by digging up a strip of uniform with along the other three sides. How wide should the border?
Answer by sofiyac(983)

A garden plot 5m by 15m has one of its sides next to a wall. The area of the plot is to be doubled by digging up a strip of uniform width along the other three sides. How wide should the border be?
---
Draw the picture of the garden with the wall on one side.
Area of the garden = 15*5 = 75 m^2
-----------------------
Let the uniform width of the stip be "x".
Area of the strip = 15x + x(x+5) + x(x+5) = 2x^2+25x
------------------------
Equation:

strip area = garden area
2x^2 + 25x = 75
2x^2+25x-75 = 0
(2x-5)(x+15) = 0
Positive solution:
x = 5/2 meters